Question 13 (1 point) A retailer uses a markup-on-selling-price approach to setting retail prices. If the retailer purchases a product from a supplier for $10 and uses a 25% markup-on- selling-price then the final retail price for the product is: $15.00 $12.50 $40.00 $13.33
Answer #1

Markup on selling price= (P-C)*100/P

25= (P-10)*100/P

25P= 100P - 1000

P= 1000/75= 13.33

Final retail price is 13.33